# Neuromorphic Overparameterisation: Generalisation and Few-Shot Learning in Multilayer Physical Neural Networks

### v4

@@ Abstract
Physical neuromorphic computing, exploiting the complex dynamics of physical systems, has seen rapid advancements in implementation sophistication and performance. Physical reservoir computing, a subset of neuromorphic computing where only the output layer is trained, faces limitations due to its reliance on single systems. This constrains output dimensionality and dynamic range, limiting performance to a narrow range of tasks and posing a significant challenge to the field’s versatility.
@@
We address this issue by engineering a suite of nanomagnetic arrays as physical reservoirs, demonstrating that tuning microstate space and geometry can significantly broaden dynamics and enhance performance. We interconnect these reservoirs in parallel and series, where each node is a distinct system to create a multilayer neural network architecture. The output of one reservoir is recorded, scaled and virtually fed as input to the next reservoir. All data is experimentally measured with outputs combined offline. This networked approach increases output dimensionality, internal dynamics and computational performance. We demonstrate that a physical neuromorphic system is capable of achieving an overparameterised state. This state facilitates meta-learning on small training sets, yielding strong performance across a wide range of tasks. Our approach’s efficacy is further demonstrated through few-shot learning, where the system rapidly adapts to new tasks.
@@
In artificial intelligence (AI) and machine learning, the performance of models often scales with their size and the number of trainable parameters. Empirical evidence supports the advantages of ’overparameterised’ regimes[1, 2] , where models, despite having a large number of parameters, avoid overfitting, generalise effectively, and learn efficiently from a limited number of training examples.
@@
Physical neuromorphic computing aims to offload processing for machine learning problems to the complex dynamics of physical systems[3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Physical computing architectures range from implementations of feed-forward neural networks[15, 16] with tuneable internal weights, to ‘reservoir’ computing, where complex internal system dynamics are leveraged for computation[8]. Neuromorphic schemes stand to benefit from the advantages of operating in an overparameterised regime. A primary use case of neuromorphic systems is in ‘edge-computing’[17] where a remotely situated device locally performs AI-like tasks. For example, an exoplanet rover vehicle performing object classification can operate more efficiently by processing data on-site rather than transmitting large datasets to cloud-servers[7]. Here, acquiring large datasets and transmitting them to cloud-servers for processing is often inefficient making the ability to compute well and adapt to new tasks with small training sets a desirable function.
@@
In physical reservoir computing, the internal dynamics of a system are not trained. Instead, only a set of weights to be applied to the readout layer are trained, reducing training costs when compared to neural network architectures. Typically, only a single physical system is employed with a fixed set of internal dynamics, resulting in a lack of versatility and overly specialised computation which is fixed at the fabrication stage. In contrast, the brain possesses a rich set of internal dynamics, incorporating multiple memory timescales to efficiently process temporal data[18]. To mimic this, research on software-based reservoirs has shown that combining multiple reservoirs with differing internal dynamics in parallel and series network architectures significantly improves performance[19, 20, 21, 22, 23, 24, 25, 26]. Parallel networks have been physically implemented[27, 28, 8], however they lack inter-node connectivity for transferring information between physical systems - limiting performance. Translating series-connected networks (often termed ‘hierarchical’ or ‘deep’) to physical systems is nontrivial and so-far unrealised due to the large number of possible inter-layer configurations and interconnect complexity.
Here, we present solutions to key outstanding problems in the physical reservoir computing field:
We fabricate three physical nanomagnetic reservoirs with high output dimensionality and show how increasing the complexity of system dynamics can improve computational properties (Nanomagnetic reservoirs, Figure 1).
We then develop and demonstrate a methodology to interconnect arbitrary reservoirs into networks. We demonstrate the computational benefits of the networked architecture and compare performance to software. Reservoir outputs are experimentally measured with network interconnections made virtually and outputs combined offline during training. (Multilayer physical neural network, Figure 2)
@@
We explore the overparameterised regime, made possible through the networking approach, where the number of network output channels far exceeds the size of the training set. The physical networks architectures do not overfit, show enhanced computational performance and are capable of fast learning with limited data. This approach is applicable to all physical systems and methods to achieve overparameterisation in an arbitrary system are discussed. (Overparameterisation, Figure 3).
We demonstrate the power of operating in an overparameterised regime by implementing a few-shot learning task using model-agnostic meta-learning[29, 30, 31]. The physical network is able to rapidly adapt to new tasks with a small number of training data points. (Learning in the overparameterised regime, Figure 4).
@@
Whilst we use nanomagnetic reservoirs to demonstrate the benefits of networking and overparameterisation, the methodology can be applied to any physical system. Additionally, we discuss the scalability of our nanomagnetic computing scheme and calculate the theoretical power consumption of a device. The scheme described here lifts the limitations of low-dimensionality and single physical systems from neuromorphic computing, moving towards a next-generation of versatile computational networks that harness the synergistic strengths of multiple physical systems. All data and code is publicly available.
@@
The physical reservoirs used here are nanomagnetic arrays which have both non-linearity and fading memory (i.e. a temporal response which depends on both current and previous inputs)[5, 32]. Each array comprises many individual nanomagnetic elements, each with its own magnetisation ‘state’. When applying an external input, the elements in an array may switch magnetisation state. The switching process depends on the external input, the current state of an element and the states of neighbouring elements through dipolar coupling, which gives rise to collective switching high-frequency dynamics which we utilise for computation. Here, we drive switching with external magnetic fields. The array state is read by measuring the absorption of injected microwaves (i.e. measuring resonance) using ferromagnetic resonance spectroscopy (FMR), producing a spectra that is highly correlated with the collective state of the array[33, 5, 34, 35].
@@
Figure 1 shows scanning electron micrographs (a,c,e) and FMR spectral evolution of three nanomagnetic reservoirs (labelled macrospin (MS), width-modifed (WM) and pinwheel (PW), see supplementary note 1 for details) when subject to a sinusoidal field-input (b,d,f). A schematic of the computing scheme is shown in Figure 1 g). The three arrays are designed to produce different and history-dependent responses to explore the effects of networking system with distinct dynamics. MS is a square lattice with bars only supporting macrospin magnetisation states. WM is a width-modified square lattice capable of hosting both macrospin and vortex spin textures. PW is a disordered pinwheel lattice with structural diversity throughout the sample. A detailed discussion of the design is provided in supplementary note 1. The high readout dimensionality of FMR is key to achieving both strong performance and overparameterisation. Other techniques for readout exist, such as magnetoresistance [36], albeit at a lower output dimensionality. During training, weights are applied to recorded data offline via ridge regression which transforms the FMR spectra to a 1D time-series with the aim of closely matching the target waveform (see Methods for further details).
@@
We assess reservoir performance via the mean-squared error (MSE) between the target and the reservoir prediction as well as two metrics: memory-capacity and nonlinearity. Memory-capacity measures the ability of the current state to recall previous inputs[37], which arises from history-dependent state evolution. Nonlinearity measures how well past inputs can be linearly mapped to the current state[37], and can arise from a number of physical system dynamics. In this work, these dynamics include resonant frequencies shifts from changing microstates and input field and the shape of FMR peaks. Metric calculations are further described in the methods section Task selection, and supplementary note 2. These metrics allow a coarse mapping between physical and computational properties, enabling comparison of different systems. Figure 1 h) shows the nonlinearity and memory-capacity metrics. Our array designs produce a diverse set of metrics, ideal for exploring the computational benefits of networking later in this work.
@@
Throughout this work, we focus on two input time-series: a sine-wave and the chaotic Mackey-Glass time-series[38] (described in the methods section Task selection). These datasets are chosen as they are compatible with 1D global field input and can be used to devise a number of computational tasks with varying requirements. We employ short training datasets (200 data points) to reflect real-world applications with strict limitations on data collection time and energy.
@@
Figure 1 i) shows the MSE of each reservoir when predicting various future values (x-axis) of the Mackey-Glass time-series. These tasks require high memory-capacity, as attaining a good prediction requires knowledge of previous inputs, with WM and PW outperforming MS due to their richer internal switching dynamics (supplementary note 1). All arrays exhibit performance breakdown at longer future steps, evidenced by the periodic MSE profiles in Figure 1 i). This is clear for the t+7 task taken from the high-MSE region (Figure 1 j), where predictions do not resemble the target. At later future steps, prediction quality improves due to the quasi-periodicity of the Mackey-Glass equation (t+11 is similar to t).
@@
Figure 1 k) shows the performance when transforming a sine-wave to cos(2x)sin(3x). The disordered PW array achieves up to 31.6 improvement for transformation tasks with no additional energy cost, highlighting the significance of system geometry and design. Performance for a history-dependent nonlinear transform task (the NARMA ‘nonlinear autoregressive moving average’ transform[40]) and further sine-transformations are shown in supplementary note 3. NARMA transforms display similar periodic profiles. For sine-transformations, no one array performs the best at all tasks, further highlighting the lack of versatility of single physical systems.
@@
To overcome the limitations of single physical systems, we now interconnect individual reservoirs to form networks, where each reservoir acts as a ‘complex node’ with memory, nonlinearity and high output-dimensionality (as opposed to traditional neural network nodes which have no memory and only 1 output dimension).
@@
We begin by making ‘parallel’ and 1D ‘series’ networks. In parallel networks, inputs and readout for each array are performed independently and combined offline (Figure 2 a). In series networks, the output from one node is used as input to the next (Figure 2 b). In this work, interconnections are made virtually, as opposed to real-time interconnection. We experimentally record all data from the first node and select one output channel to be used as input to the next node. After scaling to an appropriate field range, this time-series is then passed to the next node and its response is recorded. The full readout from both layers is then combined offline during training and prediction. We trial 49 total architectures (4 parallel and 36 2-series and 9 3-series networks with different configurations of arrays interconnections). Full details of each architecture can be found in the supplementary information. We combine the outputs of every series / parallel architecture offline to create a network that is analogous to a physical neural network (PNN) (Figure 2 c).
@@
As each node has 200 readout channels and the optimisation time required to evaluate all interconnections is not practically feasible. Seeking a more efficient solution, we explore how the output-channel characteristics from one node affect the memory-capacity and non-linearity of the of the next node in a series network. Figure 2 d) shows the per-channel non-linearity and correlation to previous inputs values for WM (see Methods for calculation details). Certain frequency channels are highly correlated with previous inputs (e.g. 7.9 - 8.1 GHz) whereas others are highly nonlinear (e.g. 7 GHz). By feeding the output of one channel as input to the next node we can evaluate how output characteristics affect the next node’s memory-capacity and non-linearity. Figure 2 e) shows this relationship when the second node is WM (other architectures shown in supplementary note 4). Corrin and NLin refer to the correlation and nonlinearity of the first node output channel. MCN2 and NLN2 are the memory-capacity and nonlinearity of the second node. Points are colour-coded depending on which array acts as the first node (consistent with previous figures). Both MCN2 and NLN2 follow an approximately linear relationship. A linear relationship is also observed when comparing correlations to specific previous inputs (supplementary note 5). As such, one may tailor the overall network metrics by selecting output-channels with certain characteristics. This interconnection control goes beyond conventional reservoir computing where interconnections are made at random[20], allowing controlled design of network properties.
@@
Figure 2 f) shows the memory-capacity and nonlinearity of selected single, parallel and series sub-networks. Parallel network nonlinearity and memory-capacity (triangular markers) lie between the nonlinearity and memory-capacity from the constituent reservoirs. Memory-capacity does not increase in parallel as no information is directly transferred between network nodes. Series networks (square and diamond symbols for series network depth of 2 and 3 respectively) show memory and non-linearity improvements above any single reservoir. The MSPWWM network (purple diamond) has both high memory and nonlinearity. PNN’s can be devised to possess any metric combination as it comprises all sub networks. The ordering of arrays in series networks is important, with memory enhancement only observed when networks are sequenced from low (first) to high (last) memory (e.g. MSWM has a larger memory capacity than WMMS), a phenomenon also seen in software reservoirs[20] and human brains[18]. To date, there has been an open question as to whether physical reservoirs are analogous to software echo-state networks. These results suggest that the two are comparable, and that methods used to improve software echo-state networks can be transferred to physical reservoir computing.
@@
We now evaluate the performance of these networks using the feature selection methodology[20, 41] described in the methods section Learning algorithms. Figure 2 g-i) compares the MSE for the best configuration of each network architecture when predicting future values of the Mackey-Glass equation (g), performing a NARMA transform (h) and combining NARMA transform and prediction (i) (additional architectures and tasks are shown in supplementary note 6). Parallel arrays (brown lines) do not show significant MSE reductions for these tasks as there is no information transfer and memory increase in this architecture. 2 and 3 series networks show significant decreases in MSE for all tasks, which improve as the series network is extended. The PNN outperforms other architectures across all future time step prediction tasks (Figure 2 g, i) with significant MSE vs flattening demonstrating higher-quality prediction. This is particularly evident when observing the predicting t+9 of the NARMA transformed Mackey-Glass equation (Figure 2 j,k) and when reconstructing the Mackey-Glass attractor (supplementary note 7). For NARMA transform (Figure 2 h), both the PNN and three-series network show a linear MSE vs profile, indicative of reaching optimum performance. Additionally, the PNN performs well across a host of nonlinear signal-transformation tasks (supplementary note 6), outperforming other network architectures at 9/20 tasks.
@@
A strength of the PNN is the massively increased readout dimensionality. The results in Figure 2 g-i) use 40 - 100 outputs for the single, parallel and series networks and 13500 outputs for the PNN. This number of outputs is far greater than the size of the training set, placing the PNN in the overparameterised regime (discussed later). In supplementary note 8, we constrain the PNN to have a similar number of outputs to the other architectures and find performance is similar to that of a three-series network. As a comparison, the grey dashed line in g-i) represents the average performance of a 50 randomly initialised 100-node echo state networks (ESN) (further software comparison provided in supplementary note 9). We find that single arrays are well matched to ESNs with 20 nodes and PNNs are well matched to ESNs with 500 nodes. Additionally, we initialise three ESNs with similar characteristics to nanomagnetic arrays and interconnect them. Interconnected ESNs, whilst functionally different, also display improvements during interconnection, highlighting both the general nature of the technique, and strengthening the analogy between physical reservoirs and software echo-state networks.
@@
We now explore the computational benefits of reaching an ‘overparameterised’ regime, where the number of network outputs exceeds the size of the training dataset. To reach overparameterisation there are three key parameters: the size of the length of the training dataset, the number of output parameters, which here refers the number of FMR channels, used during the training phase, and the effective dimensionality of the readout. The last point is critical, as the number of parameters can be arbitrarily increased by increasing the measurement resolution or repeating measurements but doing so results in increasingly correlated outputs which have no impact during training. Here, we show how engineering a rich readout response can be used to reach overparameterisation in physical systems.
@@
Figure 3 a,c) shows the train and test MSE for the best networks from each architecture (equivalent to Figure 2) when varying the number of output parameters for training set lengths of a) 50 and c) 100. b,d) Show the improvement factors () when comparing the MSEs between the underparameterised and overparameterised regime. Here, MSE is an average over 6 prediction tasks (Mackey-Glass, t+1,3,5,7,9,11) and 50 random trials of adding parameters (see methods for details).
@@
For each network and training length, we see three regimes with example predictions for each regime shown in Figure 3 e-g) for the 3 series architecture: When the number of output parameters is less than the training length, the system is underparameterised and train test MSE decrease as the number of parameters increases (Figure 3 e)). This is where the majority of neuromorphic schemes operate. When the number of output parameters is close to the length of the training set, the networks enter an overfitting regime where MSE follows a ‘U-trend’ and performance deteriorates with increasing number of outputs. Here, there are multiple solutions during training and the network is able to perfectly fit the target, leading to vanishingly small training errors. However, the fitted weights are arbitrary and when presented with unseen data, produce extreme predictions with no resemblance to the target (Figure 3 f). Surprisingly, this trend does not continue as the the number of outputs increases. Instead, the test performance substantially improves. Here, the network enters an overparameterised regime and overcomes overfitting[42, 43, 1, 2]. Instead of memorising the training data, the network can generalise and learn the underlying behaviour of the task, resulting in improved test performance (Figure 3 g). Crucially, for some networks, the overparameterised MSE is lower than the underparameterised MSE. This phenomenon has been observed in software deep learning (sometimes referred to as a double-descent phenomena[42, 43]), but not physical computing systems.
@@
The extent to which MSE recovers depends on the effective dimensionality of the readout and train length. For single reservoirs, the limited internal dynamics produce a highly correlated set of outputs. Whilst MSE reduces in the overparameterised regime, the network is unable to fully overcome the effects of overfitting resulting in high overparameterised MSE and improvement factors below 1. The networking approach increases the effective dimensionality of the readout. In parallel networks this is achieved by having distinct nodes with different dynamic responses. In series networks, nodes receive unique inputs which produce different dynamic regimes. For short training lengths (panel a,b), the enhanced output dimensionality produces a beneficial overparameterised regime with improved MSE. Improvements increase as the network size increases and even series networks with one node type can achieve a beneficial overparameterised regime (supplementary note 10). For longer train lengths (panel c,d), only the PNN reaches a beneficial overparameterised regime due to the increased effective readout dimensionality required to overcome overfitting. Task also plays a role, with more challenging prediction tasks demanding higher effective readout dimensionality to reach a beneficial overparameterised regime (supplementary note 10). Interestingly, the PNN architecture shows signatures of triple descent (Ntrain = 50, P 150 Ntrain = 100, P 600) where a peak in MSE is observed in the overparameterised regime, a sign of superabundant overparameterisation[44]. Here we use linear regression to train networks weights. When using gradient-descent, a similar trend is observed showing the results are robust to different training methods (Supplementary note 11)

### v5

# Neuromorphic Overparameterisation and Few-Shot Learning in Multilayer Physical Neural Networks

@@ Abstract
Physical neuromorphic computing, exploiting the complex dynamics of physical systems, has seen rapid advancements in sophistication and performance. Physical reservoir computing, a subset of neuromorphic computing, faces limitations due to its reliance on single systems. This constrains output dimensionality and dynamic range, limiting performance to a narrow range of tasks. Here, we engineer a suite of nanomagnetic array physical reservoirs and interconnect them in parallel and series to create a multilayer neural network architecture. The output of one reservoir is recorded, scaled and virtually fed as input to the next reservoir. This networked approach increases output dimensionality, internal dynamics and computational performance. We demonstrate that a physical neuromorphic system can achieve an overparameterised state, facilitating meta-learning on small training sets and yielding strong performance across a wide range of tasks. Our approach’s efficacy is further demonstrated through few-shot learning, where the system rapidly adapts to new tasks.
Introduction
@@
In artificial intelligence (AI) and machine learning, the performance of models often scales with their size and the number of trainable parameters. Empirical evidence supports the advantages of overparameterised regimes[1, 2] , where models, despite having a large number of parameters, avoid overfitting, generalise effectively, and learn efficiently from a limited number of training examples.
@@
Physical neuromorphic computing aims to offload processing for machine learning problems to the complex dynamics of physical systems[3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15]. Physical computing architectures range from implementations of feed-forward neural networks[15, 16] with tuneable internal weights, to reservoir computing, where complex internal system dynamics are leveraged for computation[8]. Neuromorphic schemes stand to benefit from the advantages of operating in an overparameterised regime. A primary use case of neuromorphic systems is in edge-computing[17] where a remotely situated device locally performs AI-like tasks. For example, an exoplanet rover vehicle performing object classification can operate more efficiently by processing data on-site rather than transmitting large datasets to cloud-servers[7]. Here, acquiring large datasets and transmitting them to cloud-servers for processing is often inefficient making the ability to compute well and adapt to new tasks with small training sets a desirable function.
@@
In physical reservoir computing, the internal dynamics of a system are not trained. Instead, only a set of weights to be applied to the readout layer are trained, reducing training costs when compared to neural network architectures. Typically, only a single physical system is employed with a fixed set of internal dynamics, resulting in a lack of versatility and overly specialised computation which is fixed at the fabrication stage. In contrast, the brain possesses a rich set of internal dynamics, incorporating multiple memory timescales to efficiently process temporal data[18]. To mimic this, research on software-based reservoirs has shown that combining multiple reservoirs with differing internal dynamics in parallel and series network architectures significantly improves performance[19, 20, 21, 22, 23, 24, 25, 26]. Parallel networks have been physically implemented[27, 28, 8], however they lack inter-node connectivity for transferring information between physical systems - limiting performance. Translating series-connected networks (often termed hierarchical or deep) to physical systems is nontrivial and so-far unrealised due to the large number of possible inter-layer configurations and interconnect complexity.
@@
Here, we present solutions to key problems in the physical reservoir computing field: We fabricate three physical nanomagnetic reservoirs with high output dimensionality and show how increasing the complexity of system dynamics can improve computational properties (Nanomagnetic reservoirs, Figure 1). We then develop and demonstrate a methodology to interconnect arbitrary reservoirs into networks. We demonstrate the computational benefits of the networked architecture and compare performance to software. Reservoir outputs are experimentally measured with network interconnections made virtually and outputs combined offline during training (Multilayer physical neural network, Figure 2). We explore the overparameterised regime, made possible through the networking approach, where the number of network output channels far exceeds the size of the training set. The physical networks architectures do not overfit, show enhanced computational performance and are capable of fast learning with limited data. This approach is applicable to all physical systems and methods to achieve overparameterisation in an arbitrary system are discussed (Overparameterisation, Figure 3). We demonstrate the power of operating in an overparameterised regime by implementing a few-shot learning task using model-agnostic meta-learning[29, 30, 31]. The physical network is able to rapidly adapt to new tasks with a small number of training data points (Learning in the overparameterised regime, Figure 4).
@@
Whilst we use nanomagnetic reservoirs to demonstrate the benefits of networking and overparameterisation, the methodology can be applied to any physical system. Additionally, we discuss the scalability of our nanomagnetic computing scheme and calculate the theoretical power consumption of a device. The scheme described here lifts the limitations of low-dimensionality and single physical systems from neuromorphic computing, moving towards a next-generation of versatile computational networks that harness the synergistic strengths of multiple physical systems. All data and code is publicly available[32].
Results
@@
The physical reservoirs used here are nanomagnetic arrays which have both nonlinearity and fading memory (i.e. a temporal response which depends on both current and previous inputs)[5, 33]. Each array comprises many individual nanomagnetic elements, each with its own magnetisation ‘state’. When applying an external input, the elements in an array may switch magnetisation state. The switching process depends on the external input, the current state of an element and the states of neighbouring elements through dipolar coupling, which gives rise to collective switching and high-frequency dynamics which we utilise for computation. Here, we drive switching with external magnetic fields. The array state is read by measuring the absorption of injected microwaves (i.e. measuring resonance) using ferromagnetic resonance spectroscopy (FMR), producing a spectra that is highly correlated with the collective state of the array[34, 5, 35, 36].
@@
Figure 1 shows scanning electron micrographs (Figures 1 a,c,e) and FMR spectral evolution of three nanomagnetic reservoirs (labelled macrospin (MS), width-modifed (WM)[5] and pinwheel (PW), see Supplementary Note 1 for details) when subject to a sinusoidal field-input (Figures 1 b,d,f). A schematic of the computing scheme is shown in Figure 1 g). The three arrays are designed to produce different and history-dependent responses to explore the effects of networking system with distinct dynamics. MS is a square lattice with bars only supporting macrospin magnetisation states. WM is a width-modified square lattice capable of hosting both macrospin and vortex spin textures. PW is a disordered pinwheel lattice with structural diversity throughout the sample. A detailed discussion of the design is provided in Supplementary Note 1. The high readout dimensionality of FMR is key to achieving both strong performance and overparameterisation. Other techniques for readout exist, such as magnetoresistance [37], albeit at a lower output dimensionality. During training, weights are applied to recorded data offline via ridge regression which transforms the FMR spectra to a 1D time-series with the aim of closely matching the target waveform (see Methods for further details).
@@
We assess reservoir performance via the mean-squared error (MSE) between the target and the reservoir prediction as well as two metrics: memory-capacity and nonlinearity. Memory-capacity measures the ability of the current state to recall previous inputs[38], which arises from history-dependent state evolution. Nonlinearity measures how well past inputs can be linearly mapped to the current state[38], and can arise from a number of physical system dynamics. In this work, these dynamics include resonant frequencies shifts from changing microstates and input field and the shape of FMR peaks. Metric calculations are further described in the Methods section, Task Selection, and Supplementary Note 2. These metrics allow a coarse mapping between physical and computational properties, enabling comparison of different systems. Figure 1 h) shows the nonlinearity (NL) and memory-capacity (MC) metrics. Our array designs produce a diverse set of metrics, ideal for exploring the computational benefits of networking later in this work.
@@
Throughout this work, we focus on two input time-series: a sine-wave and the chaotic Mackey-Glass time-series[39] (described in the Methods section, Task Selection). These datasets are chosen as they are compatible with 1D global field input and can be used to devise a number of computational tasks with varying requirements. We employ short training datasets (200 data points) to reflect real-world applications with strict limitations on data collection time and energy.
@@
Figure 1 i) shows the mean-squared error (MSE) of each reservoir when predicting various future values (x-axis) of the Mackey-Glass time-series. These tasks require high memory-capacity, as attaining a good prediction requires knowledge of previous inputs, with WM and PW outperforming MS due to their richer internal switching dynamics (Supplementary Note 1). All arrays exhibit performance breakdown at longer future steps, evidenced by the periodic MSE profiles in Figure 1 i). This is clear for the t+7 task taken from the high-MSE region (Figure 1 j), where predictions do not resemble the target. At later future steps, prediction quality improves due to the quasi-periodicity of the Mackey-Glass equation (t+11 is similar to t).
@@
Figure 1 k) shows the performance when transforming a sine-wave to cos(2x)sin(3x). The disordered PW array achieves up to 31.6 improvement for transformation tasks with no additional energy cost, highlighting the significance of system geometry and design. Performance for a history-dependent nonlinear transform task (nonlinear autoregressive moving average, NARMA, transform[41]) and further sine-transformations are shown in Supplementary Note 3. NARMA transforms display similar periodic profiles. For sine-transformations, no one array performs the best at all tasks, further highlighting the lack of versatility of single physical systems.
@@
To overcome the limitations of single physical systems, we now interconnect individual reservoirs to form networks, where each reservoir acts as a complex node with memory, nonlinearity and high output-dimensionality (as opposed to traditional neural network nodes which have no memory and only 1 output dimension).
@@
We begin by making parallel and 1D series networks. In parallel networks, inputs and readout for each array are performed independently and combined offline (Figure 2 a). In series networks, the output from one node is used as input to the next (Figure 2 b). In this work, interconnections are made virtually, as opposed to real-time interconnection. We experimentally record all data from the first node and select one output channel to be used as input to the next node. After scaling to an appropriate field range, this time-series is then passed to the next node and its response is recorded. The full readout from both layers is then combined offline during training and prediction. We trial 49 total architectures (4 parallel and 36 2-series and 9 3-series networks with different configurations of arrays interconnections). Full details of each architecture can be found in the supplementary information. We combine the outputs of every series / parallel architecture offline to create a network that is analogous to a physical neural network (PNN) (Figure 2 c).
@@
As each node has 200 readout channels and the optimisation time required to evaluate all interconnections is not practically feasible. Seeking a more efficient solution, we explore how the output-channel characteristics from one node affect the memory-capacity and nonlinearity of the of the next node in a series network. Figures 2 d,e) show the per-channel nonlinearity (NL) and correlation (Corr) to previous inputs values for WM (see Methods for calculation details). Certain frequency channels are highly correlated with previous inputs (e.g. 7.9 - 8.1 GHz) whereas others are highly nonlinear (e.g. 7 GHz). By feeding the output of one channel as input to the next node we can evaluate how output characteristics affect the next node’s memory-capacity and nonlinearity. Figure 2 f) shows this relationship when the second node is WM (other architectures shown in Supplementary Note 4). Corrin and NLin refer to the correlation and nonlinearity of the first node output channel. MCN2 and NLN2 are the memory-capacity and nonlinearity of the second node. Points are colour-coded depending on which array acts as the first node (consistent with previous Figures). Both MCN2 and NLN2 follow an approximately linear relationship. A linear relationship is also observed when comparing correlations to specific previous inputs (Supplementary Note 5). As such, one may tailor the overall network metrics by selecting output-channels with certain characteristics. This interconnection control goes beyond conventional reservoir computing where interconnections are made at random[20], allowing controlled design of network properties.
@@
Figure 2 g) shows the memory-capacity and nonlinearity of selected single, parallel and series sub-networks. Parallel network nonlinearity and memory-capacity (triangular markers) lie between the nonlinearity and memory-capacity from the constituent reservoirs. Memory-capacity does not increase in parallel as no information is directly transferred between network nodes. Series networks (square and diamond symbols for series network depth of 2 and 3 respectively) show memory and nonlinearity improvements above any single reservoir. The MSPWWM network (purple diamond) has both high memory and nonlinearity. PNN’s can be devised to possess any metric combination as it comprises all sub networks. The ordering of arrays in series networks is important, with memory enhancement only observed when networks are sequenced from low (first) to high (last) memory (e.g. MSWM has a larger memory capacity than WMMS), a phenomenon also seen in software reservoirs[20] and human brains[18]. To date, there has been an open question as to whether physical reservoirs are analogous to software echo-state networks. These results suggest that the two are comparable, and that methods used to improve software echo-state networks can be transferred to physical reservoir computing.
@@
We now evaluate the performance of these networks using the feature selection methodology[20, 42] described in the Methods section, Learning Algorithms. Figures 2 h-j) compares the MSE for the best configuration of each network architecture when predicting future values of the Mackey-Glass equation (h), performing a NARMA transform (i) and combining NARMA transform and prediction (j) (additional architectures and tasks are shown in Supplementary Note 6). Parallel arrays (brown lines) do not show significant MSE reductions for these tasks as there is no information transfer and memory increase in this architecture. 2 and 3 series networks show significant decreases in MSE for all tasks, which improve as the series network is extended. The PNN outperforms other architectures across all future time step prediction tasks (Figures 2 h,j) with significant MSE vs flattening demonstrating higher-quality prediction. This is particularly evident when predicting t+9 of the NARMA transformed Mackey-Glass equation (Figures 2 k,l) and when reconstructing the Mackey-Glass attractor (Supplementary Note 7). For NARMA transform (Figure 2 h), both the PNN and three-series network show a linear MSE vs profile, indicative of reaching optimum performance. Additionally, the PNN performs well across a host of nonlinear signal-transformation tasks (Supplementary Note 6), outperforming other network architectures at 9/20 tasks.
@@
A strength of the PNN is the increased readout dimensionality. The results in Figure 2 h-j) use 40 - 100 outputs for the single, parallel and series networks and 13,500 outputs for the PNN. This number of outputs is far greater than the size of the training set, placing the PNN in the overparameterised regime (discussed later). In Supplementary Note 8, we constrain the PNN to have a similar number of outputs to the other architectures and find performance is similar to that of a three-series network. As a comparison, the grey dashed line in h-j) represents the average performance of a 50 randomly initialised 100-node echo state networks (ESN) (further software comparison provided in Supplementary Note 9). We find that single arrays are well matched to ESNs with 20 nodes and PNNs are well matched to ESNs with 500 nodes. Additionally, we initialise three ESNs with similar characteristics to nanomagnetic arrays and interconnect them. Interconnected ESNs, whilst functionally different, also display improvements during interconnection, highlighting both the general nature of the technique, and strengthening the analogy between physical reservoirs and software echo-state networks.
@@
We now explore the computational benefits of reaching an overparameterised regime, where the number of network outputs exceeds the size of the training dataset. To reach overparameterisation there are three key parameters: the size of the length of the training dataset, the number of output parameters, which here refers the number of FMR channels, used during the training phase, and the effective dimensionality of the readout. The last point is critical, as the number of parameters can be arbitrarily increased by increasing the measurement resolution or repeating measurements but doing so results in increasingly correlated outputs which have no impact during training. Here, we show how engineering a rich readout response can be used to reach overparameterisation in physical systems.
@@
Figures 3 a,c) shows the train and test MSE for the best networks from each architecture (equivalent to Figure 2) when varying the number of output parameters for training set lengths of a) 50 and c) 100. Figures 3 b,d) show the improvement factors (min(MSEUP)/min(MSEOP)) when comparing the MSEs between the underparameterised (UP) and overparameterised (OP) regime. All architectures beyond single systems show improvements when operating in the overparameterised regime. Here, MSE is an average over 6 prediction tasks (Mackey-Glass, t+1,3,5,7,9,11) and 50 random trials of adding parameters (see Methods for details).
@@
For each network and training length, we see three regimes with example predictions for each regime shown in Figures 3 e-g) for the 3 series architecture: When the number of output parameters is less than the training length, the system is underparameterised and train and test MSE decrease as the number of parameters increases (Figure 3 e)). This is where the majority of neuromorphic schemes operate. When the number of output parameters is close to the length of the training set, the networks enter an overfitting regime where MSE follows a ‘U-trend’ and performance deteriorates with increasing number of outputs. Here, there are multiple solutions during training and the network is able to perfectly fit the target, leading to vanishingly small training errors. However, the fitted weights are arbitrary and when presented with unseen data, produce predictions with no resemblance to the target (Figure 3 f). Surprisingly, this trend does not continue as the the number of outputs increases. Instead, the test performance substantially improves. Here, the network enters an overparameterised regime and overcomes overfitting[43, 44, 1, 2]. Instead of memorising the training data, the network can generalise and learn the underlying behaviour of the task, resulting in improved test performance (Figure 3 g). Crucially, for some networks, the overparameterised MSE is lower than the underparameterised MSE. This phenomenon has been observed in software deep learning (sometimes referred to as a double-descent phenomena[43, 44]), but not physical computing systems.
@@
The extent to which MSE recovers depends on the effective dimensionality of the readout and train length. For single reservoirs, the limited internal dynamics produce a highly correlated set of outputs. Whilst MSE reduces in the overparameterised regime, the network is unable to fully overcome the effects of overfitting resulting in high overparameterised MSE and improvement factors below 1. The networking approach increases the effective dimensionality of the readout. In parallel networks this is achieved by having distinct nodes with different dynamic responses. In series networks, nodes receive unique inputs which produce different dynamic regimes. For short training lengths (Figures 3 a,b), the enhanced output dimensionality produces a beneficial overparameterised regime with improved MSE. Improvements increase as the network size increases and even series networks with one node type can achieve a beneficial overparameterised regime (Supplementary Note 10). For longer train lengths (Figures 3 c,d), only the PNN reaches a beneficial overparameterised regime due to the increased effective readout dimensionality required to overcome overfitting. Task also plays a role, with more challenging prediction tasks demanding higher effective readout dimensionality to reach a beneficial overparameterised regime (Supplementary Note 10). Interestingly, the PNN architecture shows signatures of triple descent (Ntrain = 50, P 150 and Ntrain = 100, P 600) where a peak in MSE is observed in the overparameterised regime, a sign of superabundant overparameterisation[45]. Here we use linear regression to train networks weights. When using gradient-descent, a similar trend is observed showing the results are robust to different training methods (Supplementary Note 11)